OpenStudy (anonymous):

Solve -2(x - 1) = 16

OpenStudy (jhannybean):

use distributive property with -2(x-1) -(2)(x) -(2)(-1) = 16 Simplify -2x +2 = 16 subtract -2 fromboth sides to keep the x'son one side and your constants on the other -2x +2 - 2 = 16 - 2 Simplify -2x = 14 divide both sides by -2 -2x/-2 = 14/-2 Simplify again x = -7

OpenStudy (anonymous):

a) 9 b) -9 c) 7 d) -7 those are my options

OpenStudy (anonymous):

-7

OpenStudy (jhannybean):

look at above post.

OpenStudy (anonymous):

jhannybean just answered you dude???

OpenStudy (jhannybean):

@Game&twerk please don't just state answers, but help a person solve the problem :)

OpenStudy (anonymous):

thank you

OpenStudy (jhannybean):

No problem.